This is about word problems, Mrs. Carey cashed a $600 check in her bank. She received a five dollar bill this is about word problems, Mrs. Carey cashed a $600 check in her bank. She received a $5 bills and $10 bills. The number of $10 bills was 10 less than three times the number of $5 bills. How many bills of each type did Mrs. Carey receive?
Answer by fractalier(6550)

Let's call the number of $5 and $10 bills, x and y.
Thus we have
y = 3x - 10 and their value
5x + 10y = 600
Now substitute
5x + 10(3x - 10) = 600
5x + 30x - 100 = 600
35x = 700
x = 20 $5 bills
y = 3(20) - 10 = 50 $10 bills